# Dissecting inertial clustering and sling dynamics in high-Reynolds number particle-laden turbulence

### v1

@@ Appendix A Backtracked trajectories
Here, we provide additional 3D visualizations of sling events in order to illustrate the back- and forward tracking that was used to produce Figs. 8 and 10. Figure 12 shows the trajectories of particles that engage in sling events at 12 $t_{c}=0.6$ (FDNS1) and 12 $t_{c}=0.4$ (FDNS5), respectively. The trajectories were obtained by tracking an ensemble of sling particles backward and forward in time from $t=t_{c}$. The color shows the fluctuation of local $R$ that is required for the sling event. Finally, we observe that at small Stokes number, the particles retain the form of a thin sheet after the sling event has taken place, whereas a pocket forms for $St_{\mathrm{eff}}=0.98$, due to higher relative velocities.
[…]

### v3

@@ Appendix A Backtracked trajectories
Here, we provide additional 3D visualizations of sling events in order to illustrate the back- and forward tracking that was used to produce Figs. 8 and 10. Figures 12(a,b) show the trajectories of particles that engage in sling events at $t_{c}=0.6$ (FDNS1) and $t_{c}=0.4$ (FDNS5), respectively. The trajectories were obtained by tracking an ensemble of sling particles backward and forward in time from $t=t_{c}$. The color shows the fluctuation of local $R$ that is required for the sling event. Finally, we observe that at small Stokes number, the particles retain the form of a thin sheet after the sling event has taken place, whereas a pocket forms for $St_{\mathrm{eff}}=0.98$, due to higher relative velocities.